OpenStudy (anonymous):

What is the value of -(-2)3?

Nnesha (nnesha):

if you take odd power of a negative number you will get negative result

OpenStudy (jdoe0001):

\(\bf -(-2)^3\implies -1\cdot (-2)^3\implies -1\cdot (-2)\cdot (-2)\cdot(-2)\implies ?\)

OpenStudy (anonymous):